1. Find the molality and molarity of the sample consisting 75% Copper and 25% Nickel by mass. The computed density of this sample is 0.00639 g/mm3.

1
Expert's answer
2021-11-17T11:32:01-0500

1 mm3 = 1x10-6 L


density = 0.00639 g / mm3 x 1x10 mm3 / L = 6390 g / L = density of the metals


mass of Cu in 1 L = 75% x 6390 g/L = 4793 g


mols Cu = 4793 g x 1 mol / 63.6 g = 75.4 mols Cu


mass of Ni in 1 L = 25% x 6390 g/L = 1598 g


mols Ni = 1598 g x 1 mol / 58.7 g = 27.2 mols Ni


volume of Cu = 0.75 L


volume of Ni = 0.25 L


volume H2O = 1 L = 0.1 L = 0.900 L


mass H2O = 0.900 kg


molality Cu = mols Cu / kg H2O = 75.4 mol / 0.90 kg = 83.8 m


molality Ni = mols Ni / kg H2O =27.2 mol / 0.90 kg = 30.2 m


molarity Cu = mols Cu / L soln = 75.4 mol / 1 L = 75.4 M


molarity Ni = mols Ni / L soln = 27.2 mol / 1 L = 27.2 M